### v3

# Fidel-TS: A High-Fidelity Multimodal Benchmark for Time Series Forecasting

Abstract
The evaluation of time series forecasting models is hindered by a critical lack of high-quality benchmarks, leading to a potential illusion of progress. Existing datasets suffer from issues ranging from pre-training data contamination in the age of LLMs to the temporal and description leakage prevalent in early multimodal designs. To address this, we formalize the core principles of high-fidelity benchmarking, focusing on data sourcing integrity, leak-free and causally sound design, and structural clarity. We introduce Fidel-TS, a new large-scale benchmark built from the ground up on these principles by sourcing data from live APIs. Our experiments reveal the flaws of the previous benchmarks and the biases in model evaluation, providing new insights into multiple existing forecasting models and LLMs across various evaluation tasks. The code is available at https://github.com/VEWOXIC/Universal-Cross-Modal-Time-Series-Forecasting-Pipeline and the data is available at https://huggingface.co/collections/wxcai/fidel-ts-68ef3c7813c4187e2c88dedd
Machine Learning, ICML
1 Introduction
The field of time series forecasting has seen relentless progress, with a proliferation of methods from classic statistical models (Box et al., 2015) and machine learning models (Chen and Guestrin, 2016) to sophisticated deep learning architectures (Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017) and foundation models (Garza et al., 2023; Rasul et al., 2023). As models demonstrate increasingly strong capabilities, the field of time-series forecasting calls for evaluations that are broader, fairer, and more aligned with real-world scenarios. However, we argue that the benchmarks currently in use have not kept pace with the rapid advancements in model development.
This weak foundation is rooted in legacy issues and compounded by new, misguided approaches. Classic unimodal benchmarks (Zhou et al., 2021; Wu et al., 2021, 2023) are often small in scale, outdated, and feature a vague composition of variables, making them scientifically unsound for evaluating modern models. More critically, the widespread availability of these datasets has created a significant, often unacknowledged, risk of pre-training contamination for large language models (LLMs) (Magar and Schwartz, 2022; Jiang et al., 2024), which may lead to the hallucination that LLMs excel in time series forecasting. As the community has pivoted to multimodal forecasting (Williams et al., 2024; Wang et al., 2024a; Xu et al., 2024a), first-generation benchmarks (Liu et al., 2024a) have perpetuated this contamination risk while introducing equally severe flaws. The common practice of retrospectively retrieving textual data introduces severe data leakage, including temporal leakage (accessing information published after the event to be predicted) and description leakage (retrieving text that explicitly states the ground-truth value). Furthermore, relying on low-frequency data (e.g., monthly) makes these datasets too small to effectively train domain-specific numerical models and prevents them from providing sufficiently complex testing scenarios.
To rectify these shortcomings and establish a more credible evaluation standard, we reconsider the essential characteristics of data encountered in real-world forecasting scenarios. Based on this review, we formalize core principles of high-fidelity benchmarking: (1) Data Sourcing Integrity, which mandates the use of real-time, authentication-protected API data streams to not only ensure recency and data sufficiency but also mitigate pre-training contamination; (2) Leak-free and Causally Sound Design, which incorporates only verifiably exogenous textual information, such as weather forecasts and scheduled maintenance, to prevent temporal and description leakage; (3) Structural Clarity, which implements a clear demarcation between forecasting "Subjects" and data "Channels" to enable rigorously evaluating model generalization. Guided by these paradigms, we introduce Fidel-TS, a new benchmark designed from the ground up to embody these principles. Each dataset contains millions of fresh, high-frequency and leak-free data points from real-time APIs, features textual information covering every timestamp, and is built with a highly extensible architecture.
To demonstrate the value of our high-fidelity benchmark, we developed a comprehensive framework compatible with models of diverse modalities and architectures. This assessment reveals performance gap compared to earlier reports, exposing the evaluation biases inherent in prior benchmarks. From this evaluation, we obtain three key findings. First, in unimodal forecasting, no model consistently dominates, and Time Series Foundation Models fail to exhibit the claimed robust zero-shot generalization. Second, multimodal gains are not inherent but depend on model architecture and the causal relevance of textual inputs. Third, although LLM performance correlates with general reasoning ability, even domain-finetuned variants remain unstable; under realistic, context-rich, and leakage-free long-term settings, LLMs suffer from severe precision and reliability issues and consistently underperform dedicated forecasting models.
In summary, our contributions are:
  • A formalization of the principles for high-fidelity time series benchmarking.
  • The release of Fidel-TS, a new large-scale benchmark built on these principles.
  • A thorough experimental analysis using our proposed framework, offering new insights into the capabilities of modern forecasting models.
2 Related Works
2.1 Unimodal Benchmarks for Time Series Forecasting
The evaluation of time series forecasting models has historically relied on a narrow set of classic unimodal benchmarks, typically including ETT(ETTh1, ETTh2, ETTm1, ETTm2), ECL (Zhou et al., 2021); Electricity, Weather, Traffic, ILI (Wu et al., 2021); M4 (Wu et al., 2023), etc. The scientific utility of these benchmarks is compromised by three core issues rooted in their age and static nature:
First, they are small in scale, with most datasets containing only on the order of tens of thousands of data points. This limited data volume is often insufficient for robust model evaluation, meaning that observed performance gains may be statistically insignificant or biased.
Second, they are outdated and have been widely accessible online for years. For instance, the ETT dataset concluded in 2018 and the Electricity dataset concluded in 2019. This long-term availability poses a significant risk of pre-training contamination, particularly for LLMs, casting doubt on the fairness of evaluations on modern LLM-based models.
Third, their variable composition is vague. There is no clear demarcation between datasets containing multiple variables from a single system and those featuring a single variable type across multiple independent systems. For example, the Weather dataset comprises various meteorological variables for atmosphere, whereas the Electricity and Traffic datasets treat sensors from different locations as distinct variables, despite them all measuring the same physical quantity.
2.2 Multimodal Benchmarks for Time Series Forecasting
While the shift to multimodality was conceptually correct, transitioning from unimodal to multimodal benchmarks is a non-trivial challenge. Classic unimodal datasets, due to their age and lack of rich contextual metadata, are difficult to extend. The initial multimodal time series forecasting benchmark, TimeMMD (Liu et al., 2024a), addressed this problem but highlighted the challenges of maintaining real world fidelity, demonstrated several drawbacks:
First, the majority of its datasets are plagued by low sampling frequencies. Six datasets are sampled monthly, including Agriculture, Climate, Economy, Security, Social Good, and Traffic, while Energy and Health are sampled weekly. This results in a severe scarcity of data points, with many series containing only a few hundred observations despite spanning decades. These simple samples cannot support the evaluation of complex scenarios or the training of numerical domain-specific models.
Second, the reliance on an Agent pipeline to retrieve web-based text introduces multiple, critical forms of data leakage. This method risks pre-training contamination, as static web content is likely already part of the training corpora for LLMs. It also leads to temporal leakage, where retrieved documents like future financial reports contain information unavailable at prediction time, and description leakage, where text explicitly describes future trends, effectively revealing the ground truth. Furthermore, this approach fails to ensure strict temporal alignment between the textual data and its corresponding time series timestamp, increasing the risk of temporal leakage issue.
Third, it perpetuates the ambiguous variable structure of unimodal benchmarks. It also fails to distinguish between datasets comprising multiple, distinct variables from a single system versus those with a single variable type from multiple systems. Besides, in its multivariate datasets, it arbitrarily designates one channel as the "target," a design choice that oversimplifies the problem and ignores the equal importance, or potential interdependencies among variables.
Table 1: Overview of Dataset Statistics in Fidel-TS
Dataset Time Span
Sampling
Frequency
No.
Subject
No. Channel
(per Subject)
Textual Data
Type
Textual Data
Frequency
Total
Data Points
Canada Calgary
Photovoltaics (CCP)
2018-01-01
$\sim$ 2025-02-28
1h 9 1 Weather 6h, 1day 564,881
Germany Renewable
Energy Grid (GREG)
2015-01-01
$\sim$ 2025-02-25
15min 8 1 Weather 6h, 1day 2,848,472
Jena Atmospheric
Physics (JAP)
2014-01-01
$\sim$ 2025-06-30
10min 1 21 Weather 6h, 1day 604,311
NYC Traffic
Speed (NYTS)
2018-01-01
$\sim$ 2025-03-01
5min 97 1 Weather 6h, 1day 67,517,110
California
ISO (CAISO)
2018-04-10
$\sim$ 2025-02-28
5min 1 20 Weather 6h, 1day 724,896
UCSD BEAR
Room (BEAR)
2020-06-01
$\sim$ 2020-10-20
5min 80 3
Weather
Control Events
6h, 1day
5min
3,156,240
Taken together, these pervasive issues of data scarcity, data leakage, and ambiguous structure across existing benchmarks result in a significant loss of real-world fidelity, further impact on the evaluation of the models.
3 Fidel-TS: A Benchmark Built on High-Fidelity Principles
Figure 1: The Construction Pipeline of Fidel-TS. The workflow integrates raw data from the Target System Time Series API and the Dynamic Heterogeneous API. The Imperfection Handling step interpolates short data gaps while converting long downtime into time-aligned ’Sensor Downtime’. These downtime events and heterogeneous textual data are stored in the Dynamic Event Database together, strictly following the start and end times. All these events and time series, combined with static dataset metadata, forms a unified dataset that is systematically organized into Realistic Benchmark Settings: downsampled (-h), sampled via importance (-mini), observed (-obs, for in-domain evaluation), and hidden (-hid, for generalization evaluation).
3.1 Core Principles of High-Fidelity Design
Our approach to achieving real-world fidelity is built on a set of interconnected principles, starting with how data is sourced and managed.
The foundation of our benchmark is the exclusive use of continuously updated data streams from authentication-protected APIs, eschewing the static, web-scraped files common in other benchmarks. This single architectural choice directly addresses several requirements for realism. First, it provides a stream of large quantity, fresh, high-frequency data, ensuring the benchmark remains current and can be used to evaluate models on fine-grained, sub-hourly tasks. This API-driven approach also guarantees precise temporal alignment across all modalities, a critical factor for accurate forecasting. Second, this methodology provides a multi-layered defense against information leakage. We mitigate the risk of pre-training contamination through two primary mechanisms. First, by sourcing data directly from APIs rather than from static, widely-indexed public files, we reduce the likelihood that our benchmark is part of common training corpora. More critically, the recency of our data streams, which are continuously updated and extend well beyond the typical knowledge cutoffs, serves as a robust safeguard against direct dataset memorization.
With a foundation of fresh, well-aligned data, our next principle is to populate it with leak-free, causally sound and context-rich textual information. We deliberately focus on weather reports (sourced from weather APIs) as the primary textual modality for the majority of our datasets. First, weather is a variable that is widely recognized in the fields of photovoltaics, electricity, and transportation as a variable that can guide the prediction rather than merely serving as a historical covariate (Antonanzas et al., 2016) (Hong and Fan, 2016) (Polson and Sokolov, 2017). Unlike economic reports or news updates, weather forecasts can be known in advance. This verifiable availability prevents temporal leakage. Second, weather data effectively circumvents description leakage. Whereas some textual inputs might explicitly describe the target variable (e.g., "power usage will rise next week"), weather reports strictly describe environmental conditions without directly stating the outcome of the variable itself. Finally, weather is causally sound and verifiable: based on real-world common sense, weather is a universal driver that significantly influences system dynamics, yet it remains an exogenous factor independent of and unaffected by the systems it influences.
To accurately reflect the structure of real-world data collection, we also formalize the distinction between subjects and channels. Our data sources, though diverse, share a common organizational structure: data originates from multiple subjects (e.g., regions, devices, or sensors), with each Subject comprising an identical set of channels (i.e., the observed variables). For example, in an electricity consumption dataset, each household is a Subject, while the measurements for ’usage (kWh)’ and ’voltage (V)’ are the Channels. A channel, therefore, refers to a distinct observed variable of a system, while a subject refers to a specific instance of that system. This clear distinction is pivotal for evaluating the generalization ability of models, as it better reflects real-world scenarios such as deploying a trained model to a newly added sensor within the same system.
Finally, the entire framework is designed for extensibility. On one hand, researchers can obtain the latest data of the dataset from the real-time API at any time through our data construction process. On the other hand, researchers can easily use the flow to add additional time series and heterogeneous data and align them in time.
3.2 Data Curation Pipeline
Translating raw, real-time data streams into a structured benchmark requires a pipeline that preserves the nuances of real-world data. The following steps detail our process, with a full summary of the complete pipeline in Figure 1 and the resulting datasets in Table 1. More detailed background information of these datasets are provided in Appendix A.1.
Handling Real-World Data Imperfections.
Embodying our principle of high fidelity, a core feature of the benchmark data is that it directly reflects the challenges of real-world instrumentation. Sourced from physical sensors rather than statistically processed tables, the raw time series data naturally contains imperfections like environmental noise and missing values. This inherent instability is a feature, not a flaw, representing the unfiltered reality of live data streams. Our first step involves filtering obvious outliers and aligning timestamps to a consistent sampling rate. To handle missing values realistically, we distinguish between short and long gaps. Short-duration gaps are filled using linear interpolation, reflecting minor data transmission issues. However, longer gaps are treated as significant real-world events; we mark the numerical values as zero and generate a corresponding textual description, such as ’sensor downtime’, which is aligned with the relevant timestamps. This transforms a data quality issue into a valuable contextual feature for the models to learn from.
Enriching Data with Static and Dynamic Textual Context.
Real-world forecasting requires both foundational knowledge and real-time updates. To provide this, we integrate two distinct types of textual information. First, we establish the static context by sourcing background knowledge from the official API metadata: a general overview for the entire dataset, specific descriptions for each Subject, and detailed explanations for each Channel. Second, we incorporate dynamic, time-aligned information that reflects the evolving state of the system and its environment. This dynamic information, along with detected sensor downtime events in Imperfection Handling step, is consolidated into the Dynamic Event Database, which features storage indexed by start and end times, ensuring that the textual context is strictly aligned with the timestamps of the numerical time series. For all datasets, we provide weather reports for their respective locations. For relevant datasets, we also support the use of control information extracted from system logs or scheduled events. See Appendix A.2 for examples and visual analysis of our preprocessed time series and textual data.
Designing Evaluation Scenarios.
To test model performance under diverse real-world conditions, we create several dataset variations. First, to ensure a fair comparison across datasets with different native sampling frequencies, we provide an option to downsample high-frequency series to a uniform hourly level (denoted by the -h suffix). Second, to address the unique computational challenges of evaluating LLMs, we curate smaller subsets (denoted by the -mini suffix) constructed via importance sampling. By selecting samples from the full test set where unimodal models outperform multimodal ones and vice versa, we ensure both computational efficiency and fairness. More details of the algorithm are provided in Appendix A.3.
Subsetting for Generalization and Transferability.
A crucial aspect of real-world utility is a model’s ability to generalize to new scenarios it has not been explicitly trained on. Our Subject and Channel definitions naturally support this evaluation. For datasets with a large number of Subjects, we create distinct partitions to simulate a zero-shot learning environment. Subsets intended for standard training and evaluation are marked with the -obs (observed) suffix, while those held back for testing generalization on unseen subjects are marked with the -hid (hidden) suffix, providing a robust test of a model’s transfer learning capabilities. More details of the division are provided in Appendix A.4.
4 Experiments
4.1 Evaluation Framework
The high-fidelity nature of our Fidel-TS benchmark establishes a robust foundation for evaluation. By supplying rich, leakage-free data, it enables not only a critical re-assessment of traditional unimodal models but also a sound basis for evaluating modern multimodal models and LLMs. However, the forecasting model landscape has become increasingly heterogeneous, with vast differences in model scale, architectural backbones, and input-output definitions. This diversity highlights a pressing need for a universal framework that enables fair and convenient comparisons.
Table 2: Comparison of time series forecasting frameworks across four key attributes
Attributes Benchmark Frameworks
LSTF- Linear TSLib TSFM- Bench MM- TSFLib Ours
Model Domain-Specific $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Foundation Models $\checkmark$ $\checkmark$
Structure Numerical $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LLM-based $\checkmark$ $\checkmark$
LLM APIs $\checkmark$ $\checkmark$
Modality Unimodal $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Multimodal $\checkmark$ $\checkmark$
Setting Full-shot $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zero-shot $\checkmark$ $\checkmark$
Existing toolkits (such as LSTF-Linear (Zeng et al., 2023), TSLib (Wang et al., 2024b), MM-TSFLib (Liu et al., 2024a), TSFM-Bench (Li et al., 2025)) often tailored to specific model categories and lack the flexibility to accommodate this wide spectrum of models under a single, unified configuration. To bridge the gap, we developed our Universal Cross-modal Evaluation Framework designed to connect our high-fidelity data with the diverse model ecosystem. The comparison in Table 2 illustrates the comprehensiveness of our benchmark across training paradigms, model types, modality support, and evaluation strategies. Our framework provides an integrated testing environment for diverse models and facilitates unified experimental configuration. More implementation details of our framework are provided in Appendix B.
4.2 Experimental Setup
Model Selection
Using our new-proposed framework, we conduct a comprehensive evaluation across representative models, which we broadly categorize into:
(1) Unimodal Forecasting Models, including MLP-based Dlinear (Zeng et al., 2023), FITS (Xu et al., 2024b), transformer-based PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024b), LLM-reprogramming-based GPT4TS (Zhou et al., 2023), and advanced foundation models Chronos (Ansari et al., 2024), TimeMoE (Shi et al., 2025) and Sundial (Liu et al., 2025), which pre-trained on vast numerical time series corpora.
(2) Multimodal Forecasting Models, including LLM-reprogramming-based GPT4MTS (Jia et al., 2024) and transformer-based FIATS (Xu et al., 2024a). They have different input requirements due to their distinct modeling approaches.
(3) LLMs, including general-purpose LLMs Qwen-2.5-14B-Instruct (Yang et al., 2024), Qwen-2.5-14B-Instruct-1M (Yang et al., 2025b), Qwen-3-14B (Yang et al., 2025a), Deepseek-R1 (DeepSeek-AI et al., 2025), and expert LLM Chattime (Wang et al., 2025) which fine-tuned based on Llama-2-7b (Touvron et al., 2023).
See Appendix C for more details of models.
Evaluation Metrics
We adopt the Mean Square Error (MSE) as the main evaluation metric:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (1)
We add a pass rate to evaluate the LLMs, defined identically to the pass@3 metric used for tasks like code generation or math problem-solving. This addresses cases where the LLM occasionally disregard prompt constraints, causing format errors (e.g., incorrect prediction length or variable numbers) and task failure.
Realistic Settings
To simulate diverse real-world time scales, our design aligns forecasting horizons with intuitive and operational periods. For instance, for datasets with an hourly frequency, we employ historical horizon of 360 (indicates half a month) and prediction lengths of {24, 168, 336, 720} which correspond to {1 day, 1 week, 2 weeks, and 1 month}. Our framework not only supports classic ratio-based splitting (e.g., 7:1:2) (Zhou et al., 2021; Wu et al., 2021), but also enables splitting by timestamps (e.g. Jan. 1st, 2021, Jan. 1st, 2022). This approach makes the evaluation setup more transparent and controllable, providing further safeguard against potential leakage issue. See Appendix D for more details of settings.
To guide our experiments, we formulate four Research Questions (RQs) and evaluate different types of models on various forecasting tasks.
4.3 RQ1: How do unimodal forecasting models perform on the new benchmark?
Table 3: Performance of unimodal forecasting models on the full test set of each dataset. All metrics are MSE. The best results are highlighted in bold. Domain-specific models are trained on the training set of each dataset. Foundation models are evaluated in zero-shot forecasting without fine-tuning on each training set.
Data -set Pred. Len. Domain-Specific Foundation Models
Patch. iTran. Dlin. FITS GPT 4TS Chro -nos Time -MoE Sun -dial
CCP -h 24 0.159 0.147 0.158 0.161 0.149 0.152 0.168 0.167
168 0.205 0.186 0.205 0.205 0.193 0.205 0.218 0.227
336 0.224 0.197 0.221 0.223 0.206 0.229 0.246 0.246
720 0.258 0.218 0.253 0.254 0.236 0.288 0.362 0.275
GREG -h 24 0.094 0.090 0.100 0.099 0.088 0.084 0.095 0.090
168 0.130 0.132 0.151 0.131 0.131 0.137 0.144 0.142
336 0.139 0.138 0.172 0.141 0.139 0.152 0.164 0.151
720 0.163 0.155 0.199 0.167 0.157 0.188 0.260 0.172
JAP -h 24 0.362 0.356 0.376 0.382 0.358 0.360 0.359 0.361
168 0.541 0.546 0.544 0.553 0.540 0.670 0.565 0.592
336 0.591 0.593 0.586 0.599 0.591 0.723 0.630 0.634
720 0.663 0.663 0.647 0.674 0.668 0.820 0.735 0.706
CAISO -h 24 0.183 0.186 0.190 0.193 0.183 0.183 0.177 0.181
168 0.360 0.386 0.392 0.368 0.366 0.389 0.377 0.382
336 0.429 0.452 0.486 0.437 0.444 0.473 0.464 0.460
720 0.536 0.555 0.655 0.550 0.542 0.608 0.610 0.577
NYTS -h-obs 24 0.601 0.581 0.638 0.644 0.582 0.613 0.680 0.649
168 1.026 1.041 1.098 1.125 1.065 1.177 1.210 1.275
336 1.176 1.195 1.262 1.306 1.194 1.512 1.421 1.497
720 1.275 1.295 1.378 1.457 1.329 1.799 1.592 1.646
BEAR -obs 12 0.121 0.105 0.128 0.169 0.113 0.151 0.147 0.141
144 0.483 0.416 0.506 0.582 0.474 1.056 0.765 0.787
288 0.680 0.619 0.650 0.767 0.652 1.312 1.012 1.043
576 0.874 0.802 0.837 1.039 0.937 1.672 1.299 1.353
To explore the ranking of the capabilities of the selected unimodal forecasting model on our benchmark, we evaluated all of them and summarized the results in Table 3. The main findings are: (1) Although most of the domain-specific models have previously claimed state-of-the-art status on classic unimodal benchmarks, the performance of them is highly dataset-dependent on our benchmark, with a different model achieving the best results on each dataset. (2) While large-scale foundation models claim to "fit all time series" and can outperform domain-specific models in zero-shot forecasting, their performance is merely comparable to specialized models on short forecast horizons and degrades significantly over longer ones. More results regarding the original sampling rate without downsampling are provided in Appendix E.1. All these suggest that the improvements observed on prior benchmarks are likely biased. On the new benchmark, no model consistently maintains a leading position, and the superior capabilities claimed by foundation models fail to materialize.
4.4 RQ2: How do models perform under the new generalization test?
Table 4: Prior and new generalization test of domain-specific unimodal forecasting models. All metrics are MSE. We report the difference in MSE between generalization tests and in-domain evaluations in Table 3; a smaller MSE gap indicates a closer match between the underlying data distributions. The best results are highlighted in bold.
Dataset Pred. Len. Patch. iTrans. Dlinear FITS GPT4TS
NYTS-h-obs $\rightarrow$ NYTS-h-hid 24 0.611 (+0.010) 0.606 (+0.025) 0.650 (+0.012) 0.655 (+0.011) 0.603 (+0.021)
168 1.123 (+0.097) 1.142 (+0.101) 1.187 (+0.089) 1.228 (+0.103) 1.162 (+0.097)
336 1.310 (+0.134) 1.329 (+0.134) 1.374 (+0.112) 1.440 (+0.134) 1.325 (+0.131)
720 1.424 (+0.149) 1.445 (+0.150) 1.501 (+0.123) 1.602 (+0.145) 1.513 (+0.184)
CCP-h $\rightarrow$ NYTS-h-hid 24 0.835 (+0.234) 0.866 (+0.285) 0.758 (+0.120) 0.763 (+0.119) 0.867 (+0.285)
168 1.308 (+0.282) 1.670 (+0.629) 1.242 (+0.144) 1.281 (+0.156) 1.334 (+0.269)
336 1.468 (+0.292) 1.928 (+0.733) 1.438 (+0.176) 1.722 (+0.416) 1.681 (+0.487)
720 1.549 (+0.274) 2.130 (+0.835) 1.568 (+0.190) 1.890 (+0.433) 5.044 (+3.715)
Prior benchmarks often employ transfer strategies that force models to transfer across datasets from entirely different domains (e.g., ETT $\rightarrow$ Weather), or across subsets within the same domain but with mismatched sampling frequencies (e.g., ETTh1 $\rightarrow$ ETTm1) (Jin et al., 2024). However, our distinction between subjects and channels enables a more realistic transfer setup. For a selected dataset, we train models exclusively on the -obs subsets and evaluate them directly on the held-out -hid subsets, simulating a more practical scenario and avoiding cross-dataset transfer.
To find the reasonable method of generalization test, we conducted both the prior and new test strategies on NYTS dataset. As shown in Table 4, the cross-dataset setting (CCP $\rightarrow$ NYTS) results in higher MSE values, with GPT4TS exhibiting an error several times larger than other baselines at length 720. Additionally, the model rankings in this setting differ from the in-domain benchmarks. In contrast, our proposed strategy (NYTS-obs $\rightarrow$ NYTS-hid) yields lower losses and maintains relative rankings that align with in-domain results in Table 3. This indicates that our method of splitting by subject and channel provides a more reasonable evaluation of generalization capability by avoiding the large distribution shifts found in cross-dataset testing. More generalization test results can be found in Appendix E.2.
4.5 RQ3: Does multimodal data actually improve predictive performance?
To verify whether multimodal forecasting models can improve performance as they claim, we conducted a comparison and summarized the results in Table 5. The results show that performance gains are not observed uniformly across all models, directly indicating that prior benchmarks suffer from bias. Furthermore, to investigate the underlying causes of the performance improvements brought by multimodal information, we leveraged our benchmark to conduct a more in-depth analysis of both models and data.
Table 5: Performance of domain-specific multimodal forecasting models on full test set of each dataset. All the models are trained on the training set of each dataset. All metrics are MSE. We report the difference between the multimodal models and the best unimodal model (as shown in Table 3), red represents better performance (MSE decrease) and blue represents worse performance (MSE increase).
Model / Dataset GPT4MTS FIATS
24 168 336 720 24 168 336 720
CCP-h 0.162 0.212 0.226 0.236 0.109 0.159 0.168 0.183
GREG-h 0.093 0.139 0.151 0.169 0.044 0.065 0.082 0.094
JAP-h 0.358 0.543 0.605 0.656 0.333 0.451 0.500 0.502
CAISO-h 0.183 0.370 0.443 0.543 0.202 0.429 0.509 0.600
NYTS-h-obs 0.617 1.099 1.290 1.335 0.407 0.644 0.685 0.710
GPT4MTS FIATS
12 144 288 576 12 144 288 576
BEAR-obs
(+Ctrl.)
0.177 0.591 0.874 1.096 0.094 0.222 0.298 0.323
BEAR-obs
(+Weather)
0.117 0.538 0.713 0.910 0.129 0.562 0.998 1.309
Model Comparison
GPT4MTS falls behind the best unimodal baseline as it restricts text usage to the historical horizon, treating it merely as a covariate that offers limited foresight. Conversely, FIATS achieves state-of-the-art performance by leveraging text within the prediction horizon, effectively using it as explicit state guidance to direct the forecasting trajectory. This difference demonstrates that the effectiveness of multimodal forecasting hinges on the model itself.
Ablation Study
The BEAR dataset provides an ideal experimental testbed for dissecting modality effectiveness, as it offers two distinct types of textual features: deterministic control events and stochastic weather reports. Results show that for the FIATS whose capabilities have already been proven by other datasets, incorporation of control events leads to a significant reduction in error, whereas the inclusion of weather data results in negligible improvements or even performance degradation. This discrepancy arises because control events are causally linked interventions that directly dictate system state changes, while weather variables exhibit a much weaker, indirect correlation with immediate system dynamics. In conclusion, the performance gains of multimodal forecasting depend on the causal relevance between textual information and the underlying time-series variables of the system.
4.6 RQ4: Are LLMs effective for forecasting tasks?
The rapid improvement in the general capabilities of LLMs has led to their evaluation on specialized time series forecasting tasks. Currently, both general and fine-tuned LLMs typically perform inference via direct prompting; for forecasting tasks, a standard approach involves constructing specific prompts based on forecasting scenario (Williams et al., 2024) (Wang et al., 2024a). To this end, we designed prompts for unimodal and multimodal contexts. Detailed prompt templates are provided in the Appendix F.1.
Table 6: Performance of LLMs on -mini test set. All metrics are MSE. The best results are highlighted in bold and the second-best are underlined. Considering the limitations of LLM context, we choose the univariate forecasting datasets and shorter prediction length (24, 168). "–" indicates that the filled-in prompt exceeds the context limit of the LLM.
Model / Setting CCP-h-mini GREG-h-mini NYTS-h-mini
24 168 24 168 24 168
Qwen2.5- 14B-Inst. 128K Uni. 0.311 0.865 0.229 0.497 1.884 3.599
Mul. 0.315 0.570 0.267 0.405 2.537 3.184
1M Uni. 0.289 0.374 0.213 0.279 1.893 2.378
Mul. 0.303 0.555 0.274 0.421 2.562 3.061
Qwen3 -14B 128K Uni. 0.235 0.298 0.153 0.263 2.236 2.546
Mul. 0.222 0.336 0.210 0.282 3.169 2.911
Deep.-R1 128K Uni. 0.182 0.267 0.113 0.213 1.071 2.154
Mul. 0.137 0.216 0.130 0.188 1.013 1.138
Chattime Uni. 0.263 0.457 0.223 0.486 0.998 2.312
Mul. 0.313 0.677 2.455
Model Comparison
The results in Table 10 yields two primary insights. First, among all evaluated LLMs, the state-of-the-art (SOTA) DeepSeek-R1 consistently achieves superior performance. The more advanced Qwen 3 also outperforms the previous generation Qwen 2.5 in most cases. This indicates that the performance ranking in forecasting tasks closely mirrors that of general-purpose capability, suggesting that robust general reasoning abilities translate effectively to time-series inference. Second, while the fine-tuned Chattime improves upon the Qwen Series and even outperforms DeepSeek-R1 in some cases, it remains unstable, particularly in multimodal scenarios. This leads to the conclusion that existing fine-tuning methods improves specific abilities but offers no stable forecasting advantage. Additional results for multivariate datasets are summarized in Appendix E.3.
Modality Comparison
For the Qwen Series and Chattime, the inclusion of multimodal information does not consistently translate into performance improvements; in many cases, the increased complexity and context length appear to burden the models, resulting in degraded predictive ability. Conversely, DeepSeek-R1 effectively leverages the additional semantic information, yielding superior results in multimodal scenarios compared to unimodal scenarios in the majority of test cases. This suggests that the utility of multimodal context is highly dependent on the model’s intrinsic reasoning capacity.
Case Study
To further investigate the source of the forecasting and reasoning capabilities of SOTA LLMs, we conducted a case study on representative examples (detailed in Appendix F.2). The study leads to the conclusion that when dealing with a unimodal scenario, DeepSeek-R1 focuses on statistical extrapolation, extracting historical trends and patterns to predict. When facing multimodal semantic context, DeepSeek-R1 use its logic-driven reasoning ability to analyze the relationship between historical weather and the time series to infer future dynamics based on provided future weather data.
Capability Concerns
Although LLMs have shown progress in forecasting tasks with the improvement of their capability, a broader comparison reveals a critical limitation: compared to the evaluation results of dedicated forecasting models on -mini subsets (in Appendix E.4), nearly all LLMs lag behind specialized baselines, while only DeepSeek-R1 surpass some of them. This suggests that previous benchmarks for evaluating the forecasting ability of LLMs may have offered overly simplified tasks with short contexts, or suffered from description leakage and temporal leakage that reduced the difficulty of prediction. In contrast, under our context-rich and leakage-free forecasting setting, LLMs encounter substantially greater challenges.
Reliability Concerns
Beside the forecasting error, we further evaluated the robustness of the LLMs given the observed failure of LLMs. The comprehensive pass@3 results are summarized in Appendix E.5. It demonstrates that while LLMs maintain high success rates in scenarios characterized by shorter prediction horizons and fewer variables, failure rates escalate substantially when the models are subjected to longer prediction length, multivariate forecasting tasks, or inputs enriched with extensive textual context. The results highlight the persistent insufficiency of current LLMs in handling long-context, long-term realistic forecasting tasks.
5 Discussion
5.1 Efficiency analysis
We summarized the parameter statistics for all models in Figure 2. For domain-specific models, linear models (FITS, DLinear) remain the most efficient, followed by transformer-based models where multimodal approaches (FIATS) surpass unimodal ones (PatchTST, iTransformer) in parameter size. For the LLM-based methods (GPT4TS/GPT4MTS), while total counts are higher, the trainable parameters remain limited due to the frozen GPT-2 backbone; The multimodal GPT4MTS is slightly larger than the unimodal GPT4TS. The parameter scale increases further with cross-domain pretrained Time Series Foundation Models (TimeMoE, Sundial, Chronos) and peaks with LLMs, ranging from SFT-based (Chattime) to the general LLMs (Qwen series and Deepseek-R1).
Figure 2: Comparison of model parameters
5.2 Limitations and Future Works
Data Perspective
To enforce a strict leak-free principle, we primarily use weather reports as textual inputs. Although weather is generally causally relevant, our ablation study shows that it is not always the most informative driver for every system. Future work will extend the dataset with more causally grounded textual sources while maintaining real-world fidelity to avoid leakage. This also calls for robust data filtering pipelines to ensure that models receive the most relevant contextual information for forecasting.
Model Perspective
We select representative models under practical computational constraints to reveal common challenges in long-term time-series forecasting. Given the rapid pace of architectural development, it is impractical to include all emerging methods, such as specific LLM prompting approaches (e.g., LLMP (Requeima et al., 2024), LLMTime (Gruver et al., 2023)). Future work can broaden these baselines to cover a wider range of models and techniques, further enhancing the long-term value of our benchmark.
6 Conclusion
In this work, we addressed a critical flaw in the evaluation of time series forecasting models by formalizing the principles of high-fidelity benchmarking and introducing Fidel-TS, a new large-scale benchmark built upon them. Our comprehensive experiments validate Fidel-TS as a rigorous paradigm, revealing a performance landscape different from prior biased benchmarks. The thorough evaluation of different categories of models reveals new insights.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
We comply with intellectual property agreements for all data sources. Data are properly anonymized and content generated by OpenAI API is free for general use, with no concerns regarding sensitive or illegal activity in our dataset.
All data sources utilized in this work are either under a CC0 license or from APIs with terms of service that permit academic research.
Appendix
Appendix A Benchmark Details
A.1 Metadata and Descriptions
Canada Calgary Photovoltaics
1. General Information: This dataset provides hourly solar power generation data, measured in kilowatt-hours (kWh), from a collection of solar panels located in Calgary, Alberta, Canada. 2. Subjects (9): A Subject in this dataset represents an individual solar panel installed at a specific public facility in Calgary. Each subject is identified by a unique ID (e.g., 314106, 319086). The metadata provides the precise location for each panel, such as "Calgary Fire Hall Headquarters" or "Southland Leisure Centre". 3. Channels (1): Each subject has a single Channel, which is the Solar Power Generation measured in kWh. 4. Textual Data: The primary source of textual information for this dataset is weather data for Calgary, Alberta.
Germany Renewable Energy Grid
1. General Information: This dataset contains time-series data on the renewable energy supply from four major transmission system operators (TSOs) in Germany. It captures the generation from both solar and wind sources. 2. Subjects (8): Each Subject represents a specific energy source from one of the German power grids. For example, the subject "solar_50Hertz" represents the solar power generation from the electricity transmission system in the north and east of Germany. 3. Channels (1): Each subject has a single Channel, representing the Energy Generation for that specific source and grid (unit is typically Megawatts, MW). 4. Textual Data: Textual information consists of regional weather data for Germany, aligned to the specific regions served by each TSO, including Bayreuth, Berlin, Dortmund, Stuttgart.
Jena Atmospheric Physics
1. General Information: This dataset consists of comprehensive atmospheric physical indicators in Jena, Germany. It is a dataset for meteorological time-series forecasting. 2. Subjects (1): There is only one Subject, which represents the entire Jena weather station as a single system. 3. Channels (21): The Channels are the 21 distinct meteorological variables measured by the station. There are some key variables such as: Carbon dioxide concentration (measured in ppm), Actual vapor pressure (measured in mbar). 4. Textual Data: The textual information is sourced from external weather forecast in Jena, Germany.
NYC Traffic Speed
1. General Information: The data is collected by traffic sensors deployed across various road segments in New York City, including information about the vehicles on the roads. 2. Subjects (97): A Subject is a unique traffic sensor monitoring a specific road segment in one of the NYC boroughs (e.g., "Queens, with segment of CIP N TNB - Whitestone Expwy S Exit 14 (Linden Pl)"). Each sensor is identified by a unique ID (e.g., 204, 184). 3. Channels (1): Each subject has a single Channel, which is the Average Vehicle Speed in km/h. 4. Textual Data: The textual data is composed of weather information for New York City, covering New York, Queens and Brooklyn.
California ISO
1. General Information: This dataset provides a holistic view of the California electricity grid, managed by the California Independent System Operator (CAISO). 2. Subjects (1): There is only one Subject, representing the entire CAISO grid system as a single, interconnected entity. 3. Channels (20): The Channels are the 20 distinct metrics that describe the state of the grid. These include demand, power generation from various sources (Hydro, Nuclear, Biomass, etc.), and CO2 emissions. 4. Textual Data: Textual information is derived from state-wide weather data for California, including San Francisco and San Diego.
UCSD BEAR Room
1. General Information: This dataset contains HVAC (Heating, Ventilation, and Air Conditioning) relevant data collected from 80 rooms within a single building (NAE-01) at the University of California San Diego. 2. Subjects (80): A Subject corresponds to an individual room within the building. Each room is identified by its unique room number (e.g., 104, 208). The metadata describes each room’s specific location. 3. Channels (3): Each subject (room) has a consistent set of three Channels: Zone Temperature, Real Power, and Supply Flow. 4. Textual Data: This dataset uniquely features two types of textual information. The primary external source is weather data for San Diego, California. Additionally, it includes internal control event logs (e.g., ’scheduled occupation of a room’).
A.2 Visualization
Figure 3: Visualization of the time series data in three datasets with a single channel. We select a representative time periods to show the patterns of the time series in each dataset.
Time Series Data
For illustrative purposes, we selected datasets with a single channel and plotted one of their subjects over a fixed time range. It can be observed from Figure 3 that the CCP and GREG datasets exhibit clear periodicity within the selected range. The NYTS dataset is not only noisier and exhibits more frequent, irregular fluctuations, leading to greater prediction difficulty, but also includes periods of sensor downtime, resulting in significant sudden distributional shifts. This suggests that predicting on the NYTS dataset is more challenging and will likely result in higher errors compared to the CCP and GREG datasets.
Figure 4: Visualization of the textual data. We select representative weather data and control events.
Textual Data
We also provide word clouds generated from one year of weather data in the CCP dataset and the control events from a single room in the BEAR dataset. As illustrated in Figure 4, these visualizations clearly highlight the main content of each text source.
A.3 Importance Sampling Method for LLM evaluation
Evaluating LLMs on the full test set presents prohibitive computational costs due to the high frequency and large scale of the data. To address this while ensuring the evaluation remains unbiased, we constructed a representative subset (denoted as -mini) using the importance sampling strategy.
We observed that samples within the test set exhibit varying degrees of multimodal dependency: for some instances, textual information is highly correlated with the time series and significantly aids forecasting, while for others, it offers marginal gains. Simple random sampling risks distorting this distribution. To preserve the original "difficulty" distribution, we employed two lightweight models: a representative unimodal model (PatchTST, (Nie et al., 2023)) and a multimodal model (FIATS, (Xu et al., 2024a)). By evaluating both on the full test set, we categorized samples into two groups based on the performance gap: (1) Text-sensitive samples: Instances where the multimodal model significantly outperformed the unimodal one. (2) Text-insensitive samples: Instances where the improvement was negligible.
Finally, we applied stratified sampling with a fixed ratio to each group. Based on the data volume of each dataset, the final selected ratios are: 1% per subject for NYTS, 5% per subject for BEAR, and 10% per subject for other datasets. This ensures an appropriate data volume for the final -mini test set, while test set faithfully reflects the multimodal sensitivity distribution of the original data, providing a reliable and efficient benchmark for LLM evaluation.
A.4 Subsetting Method for Generalization Evaluation
NYTS
The NYTS dataset comprises 97 subjects, representing traffic speed sensors distributed across New York City. Among these, 10 are newly installed sensors with data available only since August 2021. We hold out these 10 subjects to form the -hid subset, while the remaining 87 subjects with complete historical data constitute the -obs subset. This partitioning provides a high-fidelity setting that realistically simulates a common real-world scenario where new sensors are added to an existing network.
BEAR
The BEAR dataset contains complete HVAC data for 80 rooms (subjects) within a single building. We partition the subjects based on their spatial distribution. Specifically, from each floor, we randomly sample a mix of window-side and non-window-side rooms. This process yields an -obs subset of 29 subjects and a -hid subset of 10 subjects.
Appendix B Framework Implementation
Our Universal Cross-modal Evaluation Framework supports both raw text and pre-computed embedding vectors while rigorously enforcing the temporal alignment between textual data and time series. To cater to the varied nature of modern models, our framework not only provides standard PyTorch (Paszke et al., 2019) interface, but also integrates HuggingFace Transformers (Wolf et al., 2020) for foundation models, and leverages PyTorch Lightning (Falcon et al., 2019) to accelerate the training of complex multimodal models. Recognizing the unique requirements of LLMs, it also supports both local deployment via vLLM (Kwon et al., 2023) and remote API calls through a simple socket. This thoughtful engineering ensures that our framework serves as the essential bridge between the rich, multimodal data of Fidel-TS and the diverse array of models, enabling the fair and comprehensive experiments that follow.
Appendix C Model Details
Table 7 contains all the models we select for our experiments. We summarize their Training paradigm, Architecture, Modality and Textual Input Requirements. “General Info" and "Channel/Subject Info” corresponds to the dataset descriptions and the meanings of channels/subjects provided in the metadata. “History” denotes the textual information within the historical window. “Condition” refers to the conditions used for forecasting over the future horizon.
Note that LLM-based domain-specific models (GPT4TS, GPT4MTS) are distinct from LLMs. They only employ a frozen GPT-2 (Radford et al., 2004) as a backbone, and only the surrounding input and output layers are trainable. Therefore, they must be trained on the dataset and cannot perform zero-shot forecasting.
Table 7: Key Information of Models
Model Training Paradigm Architecture Modality Textual Input Requirements
PatchTST Domain-Specific Numerical Transformer Unimodal
iTransformer Domain-Specific Numerical Transformer Unimodal
Dlinear Domain-Specific Linear (MLP) Unimodal
FITS Domain-Specific Linear (MLP) Unimodal
GPT4TS Domain-Specific LLM-Reprogramming Unimodal
Chronos Cross-Domain Numerical Transformer Unimodal
TimeMoE Cross-Domain Numerical Transformer Unimodal
Sundial Cross-Domain Numerical Transformer Unimodal
GPT4MTS Domain-Specific LLM-Reprogramming Multimodal History
FIATS Domain-Specific Numerical Transformer Multimodal
Condition, General Info, Subject/Channel Info
Qwen2.5-14B-Instruct
Cross-Domain LLM
Unimodal/Multimodal
History, Condition, General Info, Subject/Channel Info
Qwen2.5-14B-Instruct-1M
Cross-Domain LLM
Unimodal/Multimodal
History, Condition, General Info, Subject/Channel Info
Qwen3-14B Cross-Domain LLM
Unimodal/Multimodal
History, Condition, General Info, Subject/Channel Info
Deepseek-R1 Cross-Domain LLM
Unimodal/Multimodal
History, Condition, General Info, Subject/Channel Info
Chattime Cross-Domain LLM
Unimodal/Multimodal
Condition, General Info, Subject/Channel Info
Appendix D Experimental Settings Details
In our experiments, we strictly follow to the principle of Realistic Settings. For all datasets except BEAR, we uniformly adopt the datasets downsampled to hourly (suffixed with -h). For these hourly datasets, we employ a historical horizon of 360 time steps (representing half a month) and prediction lengths of {24, 168, 336, 720}, corresponding to {1 day, 1 week, 2 weeks, and 1 month}, respectively.
The BEAR dataset, due to its limited time span, is not downsampled to prevent significant data loss. Consequently, we utilize its original 5-minute sampling frequency. The configuration for BEAR includes a historical horizon of 288 time steps (representing 1 day) and prediction lengths of {12, 144, 288, 576}, which correspond to {1 hour, 12 hours, 1 day, and 2 days}, respectively.
After selecting the appropriate version of each dataset, we partition them into training, validation, and testing sets. We employ the realistic split method based on timestamps. For all datasets except BEAR, we use {Jan. 1, 2021, Jan. 1, 2022}, as the split points to separate the three sets. For the BEAR dataset, the corresponding split points were set to {Sep. 1, 2020, Oct. 1, 2020}.
Specifically mentioned, for all domain-specific models, both unimodal and multimodal, training must be conducted on each dataset individually. We first train these models and subsequently test them on the full test sets. Following this, we apply our aforementioned importance sampling principle to construct the -mini test sets. For computationally expensive models such as LLMs, evaluation is performed exclusively on these lighter -mini test sets.
Appendix E Additional Experiment Results
E.1 Results of Unimodal Forecasting Models on Datasets without Downsampling
To maintain the same forecasting setting (a half-month history to predict the next day, the next week, the next 2 weeks and the next month), we adapt the history and prediction lengths to match the native frequency of each dataset. Due to the prohibitive computational cost, we selectively conducted experiments on a subset of models (for transformer-based, MLP-based, and foundation models, we select one representative model from each category) over shorter prediction horizons (day and week), with the results presented in Table 8. These results indicate that the models’ performance on the raw-resolution data is similar to that on the data downsampled to an hourly frequency.
Table 8: Performance of selected unimodal forecasting models on full test set of each dataset without downsampling. All metrics are MSE.
Dataset
Samp.
Rate
His.
Len.
Pred.
Len.
Patch. Dlinear Sundial
GREG 10min 1440 96 0.091 0.106 0.097
672 0.131 0.154 0.166
JAP 15min 2160 144 0.367 0.399 0.388
1008 0.557 0.565 0.610
CAISO 5min 4320 288 0.185 0.191 0.299
2016 0.367 0.397 0.610
NYTS-obs 5min 4320 288 0.600 0.618 0.844
2016 1.071 1.174 1.391
E.2 Generalization Test Results of BEAR Dataset
The generalization results for the BEAR dataset using our proposed strategy are presented in Table 9. Consistent with the findings on NYTS, the model rankings in this setting align with those obtained from the in-domain evaluation.
Table 9: New generalization test of domain-specific unimodal forecasting models. All metrics are MSE. The best results are highlighted in bold.
Dataset Pred. Len. Patch. iTrans. Dlinear FITS GPT4TS
BEAR-obs $\rightarrow$ BEAR-hid 12 0.100 0.059 0.105 0.139 0.094
144 0.424 0.382 0.456 0.507 0.426
288 0.588 0.549 0.595 0.659 0.579
576 0.762 0.717 0.801 0.906 0.876
E.3 Evaluation Results of LLMs on Multivariate Datasets
In multivariate forecasting scenarios, the simultaneous processing of numerous variables results in excessive token lengths, frequently causing standard 128k-context LLMs to fail due to context overflow. Consequently, we restrict our evaluation on the multivariate -mini test set to models capable of supporting extended context windows, with the performance results summarized in Table 10. Despite this filtration, we observe two distinct failure modes in these tasks. First, even when operating within theoretical limits, models often struggled to generate the full sequence, resulting in truncated outputs. Second, the overwhelming context length frequently impaired instruction adherence. Models tended to overlook the specific constraint to "predict multivariate variables while maintaining dimensions," resulting in outputs with incorrect feature dimensionality rather than the required original structure.
Table 10: Performance of LLMs on -mini test set of multivariate dataset. All metrics are MSE. Considering the limitations of LLM context, we choose the shorter prediction length (24, 168, and for BEAR is (12, 144)). Blank indicates that the filled-in prompt exceeds the input length limit of the LLM.
Model / Setting JAP -h-mini CAISO -h-mini BEAR -mini
24 168 24 168 12 144
Qwen2.5- 14B-Inst. 1M Uni. 1.825 0.501 0.522 7.425
Mul. 1.457 0.421 1.577(+Wea.) – (+Ctrl.)
Chattime Uni. 0.483 0.580 0.216 0.453 0.145 1.077
Mul. 0.624 0.874 1.209(+Wea.) – (+Ctrl)
E.4 Full Results of Forecasting Models on -mini Test Sets
To facilitate a direct comparison between LLMs and specialized forecasting models, we evaluate all baseline models on the -mini test sets, as summarized in Table 11. The performance trends and model characteristics observed on this subset align closely with those from the full test sets. This consistency attests to the fairness of our sampling algorithm, confirming that the -mini test sets effectively reduce the number of samples and thereby lower the computational overhead for LLM evaluation, while maintaining representativeness and avoiding the introduction of model-specific biases.
Table 11: Performance of all forecasting models on -mini test set of all datasets. All metrics are MSE.
Dataset Pred. Len. Domain-Specific (Unimodal) Foundation Models (Unimodal) Domain-Specific (Multimodal)
PatchTST iTrans. Dlinear FITS GPT4TS Chronos Time-MoE Sundial GPT4MTS FIATS
CCP-mini 24 0.140 0.127 0.138 0.141 0.129 0.144 0.146 0.131 0.146 0.093
168 0.192 0.169 0.197 0.197 0.179 0.203 0.219 0.192 0.199 0.142
GREG-mini 24 0.073 0.070 0.089 0.086 0.090 0.093 0.089 0.085 0.092 0.047
168 0.135 0.135 0.150 0.137 0.143 0.170 0.162 0.154 0.156 0.074
JAP-mini 24 0.286 0.285 0.334 0.334 0.278 0.277 0.278 0.286 0.275 0.234
168 0.418 0.411 0.419 0.429 0.421 0.443 0.449 0.448 0.416 0.304
NYTS-mini 24 0.826 0.764 0.864 0.871 0.774 0.880 0.923 0.777 0.838 0.387
168 1.538 1.599 1.560 1.615 1.749 1.876 1.675 1.797 1.643 0.700
CAISO-mini 24 0.136 0.140 0.143 0.145 0.137 0.146 0.132 0.130 0.138 0.170
168 0.276 0.297 0.324 0.285 0.278 0.297 0.298 0.282 0.303 0.364
BEAR-mini 12 0.103 0.080 0.111 0.144 0.091 0.126 0.118 0.116 0.093(Ctrl.) 0.102(Wea.) 0.066(Ctrl.) 0.123(Wea.)
144 0.484 0.353 0.505 0.593 0.432 0.836 0.831 1.027 0.540(Ctrl.) 0.569(Wea.) 0.211(Ctrl.) 0.546(Wea.)
E.5 Full Results of Pass@3 in LLM evaluation
To systematically evaluate the reliability of LLMs in forecasting tasks, we report the pass@3 metric (Table 12), which reveals critical stability bottlenecks tied to input complexity and context limitations. The results highlight three primary failure mechanisms:
First, the impact of High-Dimensional Data and Context Explosion. In unimodal settings, LLMs exhibit high stability on datasets with fewer variables (e.g., CCP, GREG, NYTS), maintaining pass rates above 90%. However, performance collapses on high-dimensional multivariate datasets like JAP and CAISO. Standard 128K-context models yield a 0 pass rate on these tasks. This is attributed to context explosion, where the sheer volume of historical numerical tokens from multiple channels exceeds the model’s effective window, preventing valid output generation.
Second, the "Multimodal Tax" on Reliability. Introducing textual information significantly degrades stability compared to unimodal baselines. Unlike prior works using sparse text, our benchmark involves high-frequency, dense textual data. This additional token load exacerbates the context bottleneck. For instance, on the BEAR dataset, even robust models like DeepSeek-R1 drop from a 99.9% pass rate (Unimodal) to 0 (Multimodal) as the combined input length of weather reports and historical data overwhelms the context window. Similarly, the domain-finetuned Chattime, while achieving perfect stability (100%) in unimodal tasks, fails catastrophically in multimodal long-horizon scenarios (dropping to 0.0%), indicating that current fine-tuning strategies may not sufficiently adapt models to handle such dense multimodal contexts.
Third, the prediction Horizon Sensitivity. Across all datasets and models, extending the prediction horizon (e.g., from 24 to 168) consistently lowers pass rates. This suggests that LLMs struggle to maintain coherence over long output sequences, often resulting in truncation or format deviations when generating extended forecasts. In summary, while LLMs show promise in low-resource settings, their reliability is severely compromised in realistic scenarios requiring the simultaneous processing of high-dimensional variables, long-term history, and extensive textual context.
Table 12: Summarization of all pass@3 between Unimodal (Uni.) and Multimodal (Multi.) forecasting for LLMs.
Dataset Pred. Len. Qwen2.5-14B-Instruct Qwen3-14B Chattime DeepSeek-R1
128K 1M 128K 128K
Uni. Mul. Uni. Mul. Uni. Mul. Uni. Mul. Uni. Mul.
CCP-mini 24 96.3 96.8 96.9 97.1 98.8 95.6 100.0 100.0 99.9 99.8
168 75.0 71.4 73.6 81.5 90.6 77.2 100.0 0.0 97.5 94.9
GREG-mini 24 96.9 98.1 97.6 99.4 96.7 94.3 100.0 99.2 99.8 99.8
168 74.4 91.6 73.2 86.4 88.0 77.2 100.0 0.0 93.2 97.2
JAP-mini 24 0.0 0.0 29.1 47.4 0.0 0.0 100.0 100.0 0.0 0.0
168 0.0 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0
NYTS-mini 24 93.2 95.7 93.7 97.1 97.7 96.3 100.0 99.9 99.9 100.0
168 72.9 81.8 69.4 82.4 90.0 90.0 100.0 0.0 95.3 98.2
CAISO-mini 24 0.0 0.0 45.8 69.6 0.0 0.0 100.0 48.9 0.0 0.0
168 0.0 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0
BEAR-mini 12 96.7 0.0 95.2 63.6(+Wea.) 0.0(+Ctrl.) 99.8 0.0 100.0 88.3(+Wea.) 0.0(+Ctrl.) 99.9 0.0
144 11.6 0.0 58.1 0.0 62.8 0.0 100.0 0.0 88.4 0.0
Appendix F LLM Evaluation Details
F.1 LLM Prompt Template
We provide the prompt templates used for the unimodal and multimodal forecasting tasks with LLMs.
The template for unimodal forecasting, illustrated in Figure 5 (a), incorporates only basic information about the dataset and channel, along with the formatted historical time series as input. The system prompt directly instructs the model to learn the pattern of time series and generate the forecast.
The multimodal forecasting template shown in Figure 5 (b) extends the unimodal setup by additionally providing historical textual information and future conditions in the prediction window. The corresponding system prompt directs the LLM to not only learn the pattern of time series but also focus on the association between the textual data and the time series. For both tasks, we explicitly define the output format to minimize the likelihood of formatting errors.
Notably, we design an additional prompt template for the multimodal forecasting of the BEAR dataset to accommodate its unique control events. The weather-related content is replaced with control-related information, while all other parts remained unchanged. We applied these templates only to native LLMs, including the Qwen series and Deepseek-R1. For the Chattime which fine-tuned Llama-2-7b, we utilized its original template.
Figure 5: Prompt template for LLM unimodal forecasting and multimodal forecasting
F.2 Case Study: Example of LLM Reasoning on a Forecasting Task
To illustrate how LLMs utilize known information within the context in unimodal and multimodal forecasting scenarios, we compared the reasoning processes of Deepseek-R1 under both tasks. The examples we use are derived from the CCP-mini test set. The results are shown in Figure 6.
In the unimodal forecasting scenario, the model relies exclusively on historical time-series data and implicit domain knowledge. The reasoning process focuses on identifying periodicity (daily solar cycles) and seasonal trends (increasing daylight in the Northern Hemisphere). Crucially, the model explicitly acknowledges the lack of external variables, stating, "I don’t have weather data." Consequently, it defaults to a statistical approach, assuming that recent weather will persist, and generates predictions by applying numerical multipliers derived solely from historical correlations.
In contrast, the multimodal forecasting scenario demonstrates a more robust reasoning capability by integrating textual weather data with numerical history. The model actively utilizes the provided weather forecast ("partly sunny skies") to select a specific reference day (March 10) that shares similar conditions. Instead of simple extrapolation, it employs a logic-driven adjustment: it calculates a ratio based on the difference between the current day and the reference day at a known timestamp (11 AM) and applies this scaling factor to forecast the subsequent hours. This highlights the model’s ability to ground its predictions in semantic context rather than relying solely on historical data.
Figure 6: Comparison of the Reasoning Processes of Deepseek-R1 on Unimodal and Multimodal Forecasting Tasks